# GravSphere2: A higher-order Jeans method for mass-modeling spherical stellar systems

### v2

@@ 2 Modeling stellar kinematics with GravSphere2
GravSphere is a numerical Jeans solver that was originally introduced in 2017MNRAS.471.4541R. It has since been further extended and extensively tested with mock data (2017MNRAS.471.4541R; 2018MNRAS.481..860R; 2019MNRAS.485.2010G; 2020MNRAS.498..144G; 2021MNRAS.501..978R; 2021MNRAS.505.5686C; 2022MNRAS.510.2186G; 2024MNRAS.535.1015D; 2025arXiv250303812N; 2025arXiv250418617T) including simulations deliberately chosen to explore expected bias in dynamical models due to breaking of key assumptions, such as spherical symmetry (triaxial mocks) and dynamical pseudo-equilibrium (tidally stripped mocks). A public version of this GravSphere1 code, called pyGravSphere, was introduced in 2020MNRAS.498..144G. [^3] Available at: https://github.com/AnnaGenina/pyGravSphere. Its most recent version was introduced and made publicly available in 2021MNRAS.505.5686C [^4] Available at: https://github.com/justinread/gravsphere. which improved the velocity binning with the binulator feature, improved the mass model and improved the way that the Virial Shape Parameters contributed to the likelihood. From here on, we shall refer to this version as GravSphere1.5. Julio:2023oyg also introduced and made publicly available a version of GravSphere1.5 modified to test scalar field dark matter models in dwarf galaxies. [^5] Available at: https://github.com/marianajulio/alternative
@@
A version of GravSphere1.5 with an explicit treatment of rotation has also been presented in 2025arXiv251006905J. Lastly, self-consistent inclusion of millisecond pulsar accelerations as independent kinematic tracers to constrain mass-anisotropy models in star clusters and improved photometric modeling with an $\alpha\beta\gamma$ profile were also introduced in 2025A&A...693A.104B, who used it to probe the presence of an IMBH and stellar remnant populations in Omega Centauri.
GravSphere has been a popular tool which has been used in the study of the dark matter distribution of nearby dwarf galaxies (e.g. 2018MNRAS.481..860R; 2019MNRAS.484.1401R; 2019MNRAS.485.2010G; 2020JCAP...09..004A; 2021MNRAS.501.1188A; 2021MNRAS.505.5686C; 2021arXiv211209374Z; Julio:2023oyg; 2024MNRAS.535.1015D; 2025PhRvL.134o1001Z; 2025arXiv250504475P; 2025arXiv251006905J), galaxy clusters (2023A&A...677A..24L), and to model GCs (2025A&A...693A.104B).
In this section we introduce the improved methodology of GravSphere2 for mass-modeling stellar kinematics. We will briefly summarize the key elements of previous incarnations of GravSphere, focusing here on the novel aspects with respect to these earlier versions (2017MNRAS.471.4541R; 2018MNRAS.481..860R; 2020MNRAS.498..144G; 2021MNRAS.501..978R; 2021MNRAS.505.5686C). Crucially, these improvements include the removal of binning in favor of realistic self-consistent velocity distribution functions (Sec. 2.2), the incorporation of the fourth-order Jeans equations to treat higher moments, with a more general treatment than previous approaches (Sec. 2.1), and the inclusion of higher-order PMs. Details of mass-anisotropy models are addressed in Sec. 2.3.
@@ 2.1 An extended Jeans analysis up to fourth order
GravSphere2 numerically solves the Jeans equations for a spherical, non-rotating system of tracers, with radial density profile, $\nu_{\star}(r)$, in dynamical pseudo-equilibrium (e.g. Binney2008):
@@
is the ‘velocity anisotropy’, [^6] Unlike in e.g. 2025A&A...693A.104B, $\sigma_{t}$ corresponds to the total 2D tangential velocity dispersion, rather than the 1D averaged one (thus the factor of two). This is purely notational, but we have adopted this change because we found it more convenient to work with this quantity for the subsequent parts of this section. and we assume Newtonian weak-field gravity with an enclosed mass $M(r)$ within the 3D radial coordinate $r$.
As discussed in 2025A&A...693A.104B, the assumption of non-rotation can be relaxed to some extent by solving for the second velocity moment directly, which we do. This is done by subtracting only global averages to the velocity distribution, leaving potential rotation signatures present. These are then reabsorbed when solving the Jeans equation, leading to a quadrature term that is implicitly added to the pure velocity dispersion (see e.g. the discussion in 2016ApJ...821...44F).
The solution for the radial velocity dispersion is then given by (2005MNRAS.363..705M; 1994MNRAS.270..271V):
@@
Projecting along the observable LOS and tangential and radial PM components, respectively, we obtain:
@@
By considering these three separate components, provided that one has enough data, it is possible to alleviate the well-known problem of the degeneracy between the cumulative mass, $M(r)$, and stellar velocity anisotropy, $\beta(r)$ (1990AJ.....99.1548M; Wilkinson:2001ut; 2003MNRAS.343..401L; 2009MNRAS.395...76D; 2017MNRAS.471.4541R). Further, one can resort to higher moments of the velocity distribution to place additional constraints that have also been found to alleviate this problem. One popular choice is to use VSPs (1990AJ.....99.1548M; 2014MNRAS.441.1584R; 2017MNRAS.471.4541R), which are two numbers obtained by integrating the luminosity-weighted, fourth-order LOS velocity moment to infinity, thereby marginalizing out any higher-order velocity anisotropy. This approach is used in GravSphere1 and GravSphere1.5.
The key advantage of using VSPs over explicit modeling of the higher-order moment equations is that no higher-order analog of $\beta(r)$ is required. However, a key disadvantage is that the VSPs require integrals to infinity over the surface brightness profile and fourth velocity moment, which may both be poorly constrained at large radii. Furthermore, their large radius behavior could be adversely affected by disequilibrium due to tides. Both effects can lead to bias (e.g. 2017MNRAS.471.4541R; 2018MNRAS.481..860R; 2024MNRAS.535.1015D; 2025arXiv250418617T). For this reason, in GravSphere2 we take a different approach that – as we shall show (Sec. 4) – is less biased. Instead of using VSPs, we explicitly solve the fourth-order Jeans equations, introducing a flexible functional form for the associated higher-order anisotropy that we can then marginalize over. This is more similar to the approach introduced in 2013MNRAS.432.3361R. Explicitly, we solve the fourth-order moment Jeans equations (1990AJ.....99.1548M):
@@
As shown in 2013MNRAS.432.3361R, this system of equations can be further simplified, without loss of generality, by first introducing the fourth-order anisotropy analog:
@@
which allows the first fourth-order Jeans Eq. (10) to be recast in the more familiar form:
@@
which, by comparison to the (second-order) Jeans Eq. (3), can be readily solved for the fourth-order radial velocity moment, yielding:
@@
As for the second-order counterparts, we now seek to obtain the projected quantities which are directly related to observations. To express these purely in terms of $\braket{v^{4}_{r}}$, following 2013MNRAS.432.3361R, first, it follows from Eq. (12) that:
$\frac{\partial\nu_{\star}\braket{v^{2}_{r}v^{2}_{t}}}{\partial r}=\frac{2}{3}\bigg[(1-\beta^{\prime})\frac{\partial\nu_{\star}\braket{v^{4}_{r}}}{\partial r}-\frac{\partial\beta^{\prime}}{\partial r}\nu_{\star}\braket{v^{4}_{r}}\bigg],$ (16)
then, making a substitution using the second fourth-order Jeans Eq. (10) for the first term on the right-hand side:
$\frac{\partial\nu_{\star}\braket{v^{2}_{r}v^{2}_{t}}}{\partial r}=-\frac{2}{3}(1-\beta^{\prime})\bigg[\frac{2\beta^{\prime}}{r}\nu_{\star}\braket{v_{r}^{4}}+3\nu_{\star}\sigma_{r}^{2}\frac{GM}{r^{2}}\bigg]-\frac{\partial\beta^{\prime}}{\partial r}\nu_{\star}\braket{v^{4}_{r}},$ (17)
which can then be substituted into the second fourth-order Jeans Eq. (11) to yield an expression for the tangential fourth-order moment in terms of the radial one as: [^7] It should be understood that $\beta,\beta^{\prime},M,\braket{v_{r}^{4}},\sigma_{r}$ are all, in general, functions of $r$ (we have suppressed these implicit dependencies for notational convenience).
$\braket{v^{4}_{t}}(r)=\frac{4}{3}\bigg[(1-\beta)(2-\beta^{\prime})-\frac{r}{2}\frac{\partial\beta^{\prime}}{\partial r}\bigg]\braket{v^{4}_{r}}+2(\beta-\beta^{\prime})\sigma_{r}^{2}\frac{GM}{r}.$ (18)
By considering the geometry of the problem, expanding to fourth order, and simplifying in terms of radial moments for the projected LOS component, one finds:
@@
$+\frac{3}{4}(\beta-\beta^{\prime})\sigma_{r}^{2}GM\frac{R^{4}}{r^{3}}.$ (20)
As noted in 2013MNRAS.432.3361R, this approach constitutes a significant theoretical improvement over second-order Jeans methods by ensuring physical consistency of a given mass model with higher-order velocity moments. This is not trivially warranted, as Jeans models can in some cases lead to unphysical solutions implying negative distribution functions, or be inconsistent with the velocity distribution that is implicitly assumed to map velocity dispersions (1992ApJ...391..531D; 2013MNRAS.432.3361R) (this is explicitly addressed in the Sec. 2.2). It also has the promise of serving as a means to further constrain mass models by considering additional higher-order moments (2013MNRAS.432.3361R).
One potential disadvantage, however, which is also noted in 2013MNRAS.432.3361R, is that the constraining power of this approach can be compromised due to the fact that a fully general treatment leads to the introduction of the fourth-order anisotropy analog $\beta^{\prime},$ meaning that one may not be able to unambiguously constrain solutions to the fourth-order Jeans equation for a given mass model and (second-order) anisotropy $\beta,$ leading to a problem that is similar to the original mass-anisotropy degeneracy. One approach to this problem is to impose a prescription that fixes $\beta^{\prime}.$ This can be done, for example, by assuming distribution function models which predict these (2002MNRAS.333..697L; 2005MNRAS.363..918L). More generally, as explored in 1992ApJ...391..531D; 2013MNRAS.432.3361R, assuming a particular form for the distribution function amounts to imposing a prescription to solve all the moments of the Jeans equations (beyond fourth order included). A problem with this approach, however, is that it may restrict the space of solutions unnecessarily and bias the models, unless one is able to be sufficiently judicious and general in specifying a form for the distribution function.
To tackle this potential problem, and to make full use of the available kinematic information, much like how one can use PM velocity dispersions to constrain the second-order system, we extend this methodology by also considering the fourth-order velocity moments for the tangential and radial PM components, constraining also the fourth-order system. We show that this information can be then used to further constrain the system while allowing for a general treatment of $\beta^{\prime}.$ As in Eq. (19), this leads to an expression for the tangential and radial fourth-order moments, which can be obtained (respectively) as:
@@
$+\frac{3}{4}(\beta-\beta^{\prime})\sigma_{r}^{2}\frac{GM}{r},$ (23)
@@
In this way, one can exploit the increasing abundance of PM data found in nearby galactic systems and star clusters to constrain the fourth-order anisotropy, leading to a full three-dimensional and self-consistent modeling of both the second and fourth-order moments along the three observable LOS and PM components. To the knowledge of the authors, this is a new point that has not been made in the literature before. [^8] We note that 2017ApJ...841...90E independently introduced fourth-order PMs to quantify departures from dynamical equilibrium. However, to the authors’ knowledge, our work is the first in which these are explicitly solved for from the Jeans equations and used to break the mass-anisotropy degeneracy, and is therefore a novelty in this respect.
Lastly, we emphasize the following: while PMs constitute an important and novel addition to the higher-order treatment of kinematics, this does not mean that our approach will only be useful when these are available. In particular, the fact that quantities such as VSPs exist – i.e. fourth-order anisotropy-independent constraints on the higher moments – demonstrates that the Jeans equations should automatically encode this information, yielding comparable constraints if we marginalize over some general functional forms for $\beta$ and $\beta^{\prime}$. In fact, given that virial constraints amount to positivity constraints on velocity moments and that the virial theorem can itself be derived from the Jeans equations (1992ApJ...391..531D), any quantities derived from these constraints should already be encoded in the Jeans equations. This means that the Jeans equations, solved at fourth order, should have at least as much information as using the Jeans equations at second order with VSPs. This appears to be an underappreciated fact: without additional assumptions, quantities such as VSPs should not provide any more information than velocity moments themselves from the Jeans equations (at a given order). So anisotropy-independent constraints derived from these are not inherently more constraining, but rather reflect a property of the Jeans equations themselves, which can be recovered after marginalization over the anisotropy. Therefore, one does not need to resort to these quantities in order to break the mass-anisotropy degeneracy, and we may obtain better performance by working directly with the fourth-order Jeans equations instead.
In fact, VSPs can be susceptible to significant biases, since they rely on assumptions about the behavior of the light profile and fourth velocity moment at large radii where data constraints are poor. By contrast, in the approach we take here, we avoid these potential biases but must assume instead some functional form for $\beta^{\prime}$ to marginalize over. In Sec. 4, we will demonstrate using mock data for LOS velocities without PMs that the approach we take in GravSphere2 is indeed less biased than GravSphere1 or GravSphere1.5.

### v3

@@ 2 Modeling stellar kinematics with GravSphere2
GravSphere is a numerical Jeans solver that was originally introduced in Read and Steger (2017). It has been further extended and extensively tested with mock data (Read and Steger, 2017; Read et al., 2018; Gregory et al., 2019; Genina et al., 2020; Read et al., 2021; Collins et al., 2021; Genina et al., 2022; De Leo et al., 2024; Nguyen et al., 2025; Tchiorniy and Genina, 2025), including simulations deliberately chosen to explore expected bias in dynamical models due to breaking of key assumptions, such as spherical symmetry (triaxial mocks) and dynamical pseudo-equilibrium (tidally stripped mocks). A public version of this GravSphere1 code, called pyGravSphere, was introduced in Genina et al. (2020). [^4] Available at: https://github.com/AnnaGenina/pyGravSphere. Its most recent version was introduced and made publicly available in Collins et al. (2021) [^5] Available at: https://github.com/justinread/gravsphere.. It improved the velocity binning with the binulator feature, improved the mass model and improved the way that the VSPs contributed to the likelihood. Henceforth, we refer to this version as GravSphere1.5. Júlio et al. (2023) also introduced and made publicly available a version of GravSphere1.5 modified to test scalar field dark matter models in dwarf galaxies. [^6] Available at: https://github.com/marianajulio/alternative
@@
A version of GravSphere1.5 with an explicit treatment of rotation has also been presented in Júlio et al. (2025). Lastly, self-consistent inclusion of millisecond pulsar accelerations as independent kinematic tracers to constrain mass-anisotropy models in star clusters and improved photometric modeling with an $\alpha\beta\gamma$ profile were also introduced in Bañares-Hernández et al. (2025), who used it to probe the presence of an IMBH and stellar remnant populations in Omega Centauri.
GravSphere has been a popular tool which has been used in the study of the dark matter distribution of nearby dwarf galaxies (e.g., Read et al. 2018, 2019; Gregory et al. 2019; Alvarez et al. 2020; Alvey et al. 2021; Collins et al. 2021; Zoutendijk et al. 2021; Júlio et al. 2023; De Leo et al. 2024; Zimmermann et al. 2025; Pickett et al. 2025; Júlio et al. 2025), galaxy clusters (Li et al., 2023), and to model GCs (Bañares-Hernández et al., 2025). In this section, we introduce the improved methodology of GravSphere2 for the mass modeling of stellar kinematics. We briefly summarize the key elements of previous incarnations of GravSphere, focusing here on the novel aspects with respect to these earlier versions (Read and Steger, 2017; Read et al., 2018; Genina et al., 2020; Read et al., 2021; Collins et al., 2021). Crucially, these improvements include the removal of binning in favor of realistic self-consistent velocity distribution functions (Sect. 2.2), the incorporation of the fourth-order Jeans equations to treat higher moments, with a more general treatment than previous approaches (Sect. 2.1), and the inclusion of higher order PMs. The details of mass-anisotropy models are addressed in Sect. 2.3.
@@ 2.1 An extended Jeans analysis up to fourth order
GravSphere2 numerically solves the Jeans equations for a spherical, nonrotating system of tracers, with radial density profile, $\nu_{\star}(r)$, in a dynamical pseudo-equilibrium (e.g., Binney and Tremaine, 2008), expressed as
@@
is the velocity anisotropy [^7] Unlike in, e.g., Bañares-Hernández et al. (2025), $\sigma_{t}$ corresponds to the total 2D tangential velocity dispersion, rather than the 1D averaged one (thus the factor of two). This is purely notational, but we adopted this change because we found it more convenient to work with this quantity for the subsequent parts of this section. and we assume Newtonian weak-field gravity with an enclosed mass, $M(r)$, within the 3D radial coordinate, $r$.
As discussed in Bañares-Hernández et al. (2025), the assumption of nonrotation can be relaxed to some extent by solving for the second velocity moment directly. This is what we aim to do in this study. This is done by subtracting only global averages to the velocity distribution, leaving potential rotation signatures present. These are then reabsorbed when solving the Jeans equation, leading to a quadrature term that is implicitly added to the pure velocity dispersion (see, e.g., the discussion in Fritz et al., 2016).
The solution for the radial velocity dispersion (Mamon and Łokas, 2005; van der Marel, 1994) is then given by
@@
Projecting along the observable LOS and tangential and radial PM components, respectively, we obtain
@@
By considering these three separate components, provided that one has enough data, it is possible to alleviate the well-known problem of the degeneracy between the cumulative mass, $M(r)$, and stellar velocity anisotropy, $\beta(r)$ (Merrifield and Kent, 1990; Wilkinson et al., 2002; Łokas and Mamon, 2003; de Lorenzi et al., 2009; Read and Steger, 2017). Further, we can resort to higher moments of the velocity distribution to place additional constraints that have also been found to alleviate this problem. One popular choice is to use VSPs (Merrifield and Kent, 1990; Richardson and Fairbairn, 2014; Read and Steger, 2017), which are two numbers obtained by integrating the luminosity-weighted, fourth-order LOS velocity moment to infinity, thereby marginalizing out any higher order velocity anisotropy. This approach is used in GravSphere1 and GravSphere1.5.
The key advantage of using VSPs over explicit modeling of the higher order moment equations is that no higher order analog of $\beta(r)$ is required. However, a key disadvantage is that the VSPs require integrals to infinity over the surface brightness profile and fourth velocity moment, which may both be poorly constrained at large radii. Furthermore, their large radius behavior could be adversely affected by disequilibrium due to tides. Both effects can lead to bias (e.g., Read and Steger, 2017; Read et al., 2018; De Leo et al., 2024; Tchiorniy and Genina, 2025). For this reason, in GravSphere2 we take a different approach that (as we show in Sect. 4) is less biased. Instead of using VSPs, we explicitly solve the fourth-order Jeans equations, introducing a flexible functional form for the associated higher order anisotropy that we can then marginalize over. This is more similar to the approach introduced in Richardson and Fairbairn (2013). Explicitly, we solve the fourth-order moment Jeans equations (Merrifield and Kent, 1990), expressed as
@@
As shown in Richardson and Fairbairn (2013), this system of equations can be further simplified, without loss of generality by first introducing the fourth-order anisotropy analog,
@@
which allows the first fourth-order Jeans expression in Eq. (10) to be recast in the more familiar form of
@@
which, by comparison to the (second-order) Jeans Eq. (3), can be readily solved for the fourth-order radial velocity moment, yielding
@@
As for the second-order counterparts, we now want to obtain the projected quantities that are directly related to observations. To express these purely in terms of $\braket{v^{4}_{r}}$, following Richardson and Fairbairn (2013), first, it follows from Eq. (12) that
$\frac{\partial\nu_{\star}\braket{v^{2}_{r}v^{2}_{t}}}{\partial r}=\frac{2}{3}\bigg[(1-\beta^{\prime})\frac{\partial\nu_{\star}\braket{v^{4}_{r}}}{\partial r}-\frac{\partial\beta^{\prime}}{\partial r}\nu_{\star}\braket{v^{4}_{r}}\bigg].$ (16)
Next, we made a substitution using the second fourth-order Jeans Eq. (10) for the first term on the right-hand side,
$\frac{\partial\nu_{\star}\braket{v^{2}_{r}v^{2}_{t}}}{\partial r}=-\frac{2}{3}(1-\beta^{\prime})\bigg[\frac{2\beta^{\prime}}{r}\nu_{\star}\braket{v_{r}^{4}}+3\nu_{\star}\sigma_{r}^{2}\frac{GM}{r^{2}}\bigg]-\frac{2}{3}\frac{\partial\beta^{\prime}}{\partial r}\nu_{\star}\braket{v^{4}_{r}},$ (17)
which can then be substituted into the second fourth-order Jeans Eq. (11) to yield an expression for the tangential fourth-order moment in terms of the radial one as [^8] It should be understood that $\beta,\beta^{\prime},M,\braket{v_{r}^{4}},\sigma_{r}$ are all, in general, functions of $r$ (we suppressed these implicit dependencies for notational convenience).
$\braket{v^{4}_{t}}(r)=\frac{4}{3}\bigg[(1-\beta^{\prime})(2-\beta^{\prime})-\frac{r}{2}\frac{\partial\beta^{\prime}}{\partial r}\bigg]\braket{v^{4}_{r}}+2(\beta^{\prime}-\beta)\sigma_{r}^{2}\frac{GM}{r}.$ (18)
By considering the geometry of the problem, expanding to fourth order, and simplifying in terms of radial moments for the projected LOS component, we find
@@
$+\frac{3}{4}(\beta^{\prime}-\beta)\sigma_{r}^{2}GM\frac{R^{4}}{r^{5}}.$ (20)
As noted in Richardson and Fairbairn (2013), this approach constitutes a significant theoretical improvement over second-order Jeans methods by ensuring physical consistency of a given mass model with higher order velocity moments. This is not trivially warranted, as Jeans models can in some cases lead to unphysical solutions implying negative distribution functions or be inconsistent with the velocity distribution that is implicitly assumed to map velocity dispersions (Dejonghe and Merritt, 1992; Richardson and Fairbairn, 2013). We note that this is explicitly addressed in Sect. 2.2. It also has the promise of serving as a means to further constrain mass models by considering additional higher order moments (Richardson and Fairbairn, 2013).
One potential disadvantage, however, which was also noted in Richardson and Fairbairn (2013), is that the constraining power of this approach can be compromised due to the fact that a fully general treatment leads to the introduction of the fourth-order anisotropy analog, $\beta^{\prime},$ meaning that one may not be able to unambiguously constrain solutions to the fourth-order Jeans equation for a given mass model and (second-order) anisotropy, $\beta,$ leading to a problem that is similar to the original mass-anisotropy degeneracy. One approach to this problem is to impose a prescription that fixes $\beta^{\prime}.$ This can be done, for example, by assuming distribution function models which predict these (Łokas, 2002; Łokas et al., 2005). More generally, as explored in Dejonghe and Merritt (1992); Richardson and Fairbairn (2013), assuming a particular form for the distribution function amounts to imposing a prescription to solve all the moments of the Jeans equations (beyond the fourth order included). One problem with this approach, however, is that it might restrict the space of solutions unnecessarily and bias the models, unless one is able to be sufficiently judicious and general in specifying a form for the distribution function.
To tackle this potential problem and to make full use of the available kinematic information, similarly to how we can use PM velocity dispersions to constrain the second-order system, we extend this methodology by also considering the fourth-order velocity moments for the tangential and radial PM components, constraining also the fourth-order system. We show that this information can be then used to further constrain the system while allowing for a general treatment of $\beta^{\prime}.$ As in Eq. (19), this leads to an expression for the tangential and radial fourth-order moments, which can be obtained (respectively) as
@@
$+\frac{3}{4}(\beta^{\prime}-\beta)\sigma_{r}^{2}\frac{GM}{r},$ (23)
@@
In this way, we can exploit the increasing abundance of PM data found in nearby galactic systems and star clusters to constrain the fourth-order anisotropy, leading to a full three-dimensional and self-consistent modeling of both the second and fourth-order moments along the three observable LOS and PM components. To our knowledge, this is a new point that has not been published in the literature to date. [^9] We note that Evslin and Del Popolo (2017) independently introduced fourth-order PMs to quantify departures from dynamical equilibrium. However, to the authors’ knowledge, our work is the first in which these are explicitly solved for from the Jeans equations and used to break the mass-anisotropy degeneracy and it is therefore a novelty in this respect.
Lastly, we emphasize the following point: while PMs constitute an important and novel addition to the higher order treatment of kinematics, this does not mean that our approach will only be useful when these are available. In particular, the fact that quantities such as VSPs exist (i.e., fourth-order anisotropy-independent constraints on the higher moments) demonstrates that the Jeans equations should automatically encode this information, yielding comparable constraints if we marginalize over some general functional forms for $\beta$ and $\beta^{\prime}$. In fact, given that virial constraints amount to positivity constraints on velocity moments and that the virial theorem can itself be derived from the Jeans equations (Dejonghe and Merritt, 1992), any quantities derived from these constraints should already be encoded in the Jeans equations. This means that the Jeans equations, solved at fourth order, should have at least as much information as using the Jeans equations at second order with VSPs. This appears to be an underappreciated fact: without additional assumptions, quantities such as VSPs should not provide any more information than velocity moments themselves from the Jeans equations (at a given order). Thus, any anisotropy-independent constraints derived from these are not inherently more constraining, but rather reflect a property of the Jeans equations themselves, which can be recovered after marginalization over the anisotropy. Therefore, we do not need to resort to these quantities to break the mass-anisotropy degeneracy and it may be possible to obtain better performance by working directly with the fourth-order Jeans equations instead.
In fact, VSPs can be susceptible to significant biases, since they rely on assumptions about the behavior of the light profile and fourth velocity moment at large radii where data constraints are poor. By contrast, in the approach we take here, we managed to avoid these potential biases; however, we had to instead assume some functional form for $\beta^{\prime}$ to marginalize over. In Sect. 4, we use mock data for the LOS velocities without PMs to demonstrate that the approach we take in GravSphere2 is indeed less biased than GravSphere1 or GravSphere1.5.